Balance Sheet Analysis Complete the balance sheet and sales information using the following financial data: Total assets turnover: 1.2x Days sales outstanding: 35.5 daysa Inventory turnover ratio: 5x Fixed assets turnover: 3x Current ratio: 1.8x Gross profit margin on sales: (Sales-Cost of goods sold)/Sales-2596 Calculation is based on a 365-day year. Round your answer to the nearest cent. Balance Sheet Cash Accounts receivable Inventories Fixed assets Current liabilities $ Long-term debt 33,750 Common stock Retained earnings Total liabilities and equity Cost of goods sold 67,500 Total assets $225,000 Sales
Answer #1

Assets Turnover = Sales / Total Assets = 1.2x

Total Assets figure is given = 225,000

Hence, Sales = 1.2 * 225,000 = 270,000

Days sales outstanding = 365 / (Sales / Accounts recievables )

35.5 = 365 / Accounts recievables turnover

Accounts recievables turnover = 365 / 35.5 = 10.28x

Accounts recievables turnover = (Sales / Accounts recievables)

10.28 = 270,000 / Accounts recievables

26,264.59 = Accounts recievables

Gross profit margin = 25%

25% = Gross Profit / Revenue

25% = Gross Profit / 270,0000

25% * 270,000 = 67,500 = Gross Profit

Hence Cost of Goods Sold (COGS) = 270,000 - 67,500 = 202,500

Inventory turnover ratio = 5x

Inventory turnover ratio = COGS / Inventory

5 = 202,500 / Inventory

Inventory = 202,500 / 5 = 40,500

Fixed assets turnover = Revenue / Fixed assets

3 = 270,000 / Fixed assets

Fixed assets = 90,000

Cash = Total Assets - Accounts recievables - Inventory - Fixed assets

Cash = 225,000 - 26,264.59 -  40,500 - 90,000 = 68,235.41

Current Assets = Accounts recievables + Inventory + Cash

Current Assets = 26,264.59 + 40,500 +  68,235.41 = 135,000

Current Ratio = Current Assets / Current Liabilities

1.8 = 135,000 / Current Liabilities

Current Liabilities = 75,000

Total liabilities and equity = Total assets

Hence, Total liabilities and equity = 225,000

Total Liabilities = Current Liabilities + Long term debt

Total Liabilities = 75,000 + 33,750 = 108,750

Total Equity = Total liabilities and equity - Total Liabilities

Total Equity = 225,000 - 108,750 = 116,250

Common Equity = Total equity - retained earnings

Common Equity = 116,250 - 67,500 = 48,750